We have the survey data on the body mass index (BMI) of 642 young women. The mean BMI in the sample was x¯=26.1. We treated these data as an SRS from a Normally distributed population with standard deviation ?=7.3.

Give confidence intervals for the mean BMI and the margins of error for 90%, 95%, and 99% confidence.

Conf. Level Interval (±±0.01) margins of error (±±0.0001)
90% to
95% to
99% to
Answer #1

Solution :

Given that,

\bar x = 26.1

\sigma = 7.3

n = 642

a ) At 90% confidence level the z is ,

\alpha  = 1 - 90% = 1 - 0.90 = 0.1

\alpha / 2 = 0.1 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 1.645 * (7.3 / \sqrt642) = 0.47

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

26.1 - 0.47 < \mu < 26.1 + 0.47

25.63< \mu < 26.57

(25.63 , 26.57)

b ) At 95% confidence level the z is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 1.960 * (7.3/ \sqrt642) = 0.56

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

26.1 - 0.56 < \mu < 26.1 + 0.56

25.54 < \mu < 26.66

(25.54, 26.66)

c )At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 2.576 * (7.3 / \sqrt642) = 0.74

At 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

26.1 - 0.74 < \mu < 26.1 + 0.74

25.36 < \mu < 26.84

(25.36, 26.84)